OpenStudy (anonymous):

find the intercepts 2x+5y=10

OpenStudy (anonymous):

to find the intercept of x ,substitute 0 in the place of y 2x+0=10 2x =10 x=5(the x-intercept=5) similarly , 2(0)+5y=10 5y=10 y=2(y-intercept)